OpenStudy (anonymous):

2/5 divided by (- 7/10)* (-7/6)

OpenStudy (saifoo.khan):

2/3

OpenStudy (anonymous):

is that simplifiled?

OpenStudy (saifoo.khan):

yup!

OpenStudy (anonymous):

what? no it is 24/49

OpenStudy (saifoo.khan):

this is your prob right?

OpenStudy (anonymous):

but that depends on parenthesis I calculated (2/5) / ((- 7/10)* (-7/6))

jhonyy9 (jhonyy9):

this want to be 2/5/49/60 ?

jhonyy9 (jhonyy9):

- so if yes than you get ( 2/5) x (60/49) = 120/245 =24/49